OpenStudy (anonymous):

Form a polynomial f(x) with real coefficients having given degree and zeros. degree 5; zeros -8;-i;4+i

OpenStudy (anonymous):

Well a polynomial, by defintion of the linear factorization theorm states that a degree of n>1, then it can be written in linear factors with its zeros. It is written in the form (x-a)(x-b) where this is a 2nd degree polynomial with zeros a and b If x = a is a zero, then (x - a) is a factor. If a + bi is a zero, then a - bi is a zero as well. So see how you have 3 zeros given but this is a 5th degree polynomial? That neans it has 5 factors, but where are the other two? Well they are the cnjugates of -i and 4+i, do you know what those are?

OpenStudy (anonymous):

umm I really don't know.... I have spent over 2hours trying to figure this problem out

OpenStudy (anonymous):

Please Help!!!!!

OpenStudy (anonymous):

Lol, ok we can work through this and I'll stay until you get it, do you know how you have real numbers, numbers you can write like 1, 2, 3.14159..., 8.99579, 22/7, and then you have imaginary numbers such as \[\Large i=\sqrt{-1}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well you can sometimes write them together to make what is called a "complex number" it it written in the form \[\Large a+bi\] where a and b are real numbers and i is the imaginary number

OpenStudy (anonymous):

really any number can be written as a complex number, like \[\Large 1+2i\] is a complex number but so is 2, 2 is a complex number because it can be written like this \[\Large 2+0i\] becasue 0 times anything is still 0, and -3i can be written as a complex number because its the same thing as \[\Large 0-3i\]

OpenStudy (anonymous):

OK, sorry im back anyways, a complex number ahs a conjugate and the conjugate can be described as saying if you have a complex number (rmemeber any number real or imaginary or a combo of both can be written as this way) in the form \[\Large a+bi\] then its conjuagte can be denoted as \[\Huge a-bi\]

OpenStudy (anonymous):

Now, theres a theorem called the fundamental theormrem of algebra that states that if you have a polynomial with n-degree, it has n-roots (zeros) so in this case its a 5 degree, so it has 5 zeros. Also it says that if a complex number and its conjugate is also a zero so if they say -i is a zero then its conjugate ( this can be written as 0-i=0+i=i) is i. so i and -i are zeros tot his function. Also 4+i is zero meaning that 4-i is also a zero so now that accounts for two more zeros so now you can wirte it in linear factorization form \[\Large (x-(-8)\times(x-(i))\times(x-(-i))\times(x-(4+i))\times(x-(4-i))\]

OpenStudy (anonymous):

wow, is the last number (x-6)

OpenStudy (anonymous):

no sorry it flew off the page \[ (x-(-8)\times(x-(i))\times(x-(-i))\times(x-(4+i))\times(x-(4-i))\]

OpenStudy (anonymous):

:) thanks sooooo much... your the best of the class this year!!!!lol

OpenStudy (anonymous):

Did you understand my explanaation? Also all you have to do now is multiply the two last ones together then the two ones aftert hat and then multiply by (x-(-8))=(x+8) btw

OpenStudy (anonymous):

yes I did, and plus I took notes.. thanx!!!

OpenStudy (anonymous):

np, thats what we're here for :)